# Integrable open elliptic Toda chain with boundaries

### v1

@@ 6 Open elliptic Toda chain with boundaries
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\log\Big(2\nu_{0}^{+}\cosh({\bf{p}}_{1}/2c)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)+}\\ \\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\log\Big(2\nu_{0}^{-}\cosh({\bf{p}}_{n}/2c)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)+}\\ \\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (39)
@@
Example 1: pure open chain. Let $\nu^{\pm}_{k}=0$, $k=1,2,3$ and $\nu^{\pm}_{0}=1$. Then we come to a ”pure” open elliptic Toda chain (corresponding to $K^{\pm}(z)=\sigma_{0}$ in XYZ chain):
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{pure\ open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)\cosh({\bf{p}}_{1}/2c)}+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)\cosh({\bf{p}}_{n}/2c)}+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (40)
Example 2: boundary terms. Let $\nu^{\pm}_{0}=0$. Then
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{boundary}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp(q_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp(q_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (41)
In conclusion, we note again that the Ruijsenaars-Toda chain [2] generalizes the elliptic Toda chain (1), and it is gauge equivalent to a generic XYZ chain (without restrictions (17)). Similarly to the results of the present paper it is then possible to derive the open Ruijsenaars-Toda chain with boundaries. The boundary $K$-matrices can be also chosen to be dynamical, see the Zhukovsky-Volterra gyrostat [7] and the 4-constant ${\rm BC}_{1}$ Ruijsenaars-van Diejen model [8]. These generalizations will be studied elsewhere.
@@ Acknowledgments.
This work was supported by the Russian Science Foundation under grant no. 25-11-00081, https://rscf.ru/project/25-11-00081/ .

### v3

@@ 6 Open elliptic Toda chain with boundaries
$\begin{array}[]{c}\displaystyle{\big(b_{1}^{[2,1]}\,,b_{2}^{[2,1]}\big){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=}\\ \\ \displaystyle{=\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\big(e^{{\bf{p}}_{1}/2c}\,,-e^{-{\bf{p}}_{1}/2c}\big){\tilde{K}}^{+}(0)\left(\begin{array}[]{c}e^{{\bf{p}}_{1}/2c}\\ e^{-{\bf{p}}_{1}/2c}\end{array}\right)=}\\ \\ \displaystyle{=\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\Big(2\nu_{0}^{+}\sinh\big(\frac{{\bf{p}}_{1}}{c}\big)-4\sinh^{2}\big(\frac{{\bf{p}}_{1}}{2c}\big){\tilde{f}}^{+}({\bf{q}}_{1})\Big)=}\\ \\ \displaystyle{=4\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\Big(\nu_{0}^{+}\cosh\big(\frac{{\bf{p}}_{1}}{2c}\big)-\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big){\tilde{f}}^{+}({\bf{q}}_{1})\Big)=}\\ \\ \displaystyle{=4\sinh^{2}\big(\frac{{\bf{p}}_{1}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)\,.}\end{array}$ (72)
Similarly,
$\begin{array}[]{c}\displaystyle{\big(b_{1}^{[n-1,n]}\,,b_{2}^{[n-1,n]}\big){\tilde{K}}^{-}(0)e^{P^{n}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=}\\ \\ \displaystyle{=4\sinh\big(\frac{{\bf{p}}_{n}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{n}-{\bf{q}}_{n-1})\theta_{1}({\bf{q}}_{n}+{\bf{q}}_{n-1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{n})}\Big(\nu_{0}^{-}\cosh\big(\frac{{\bf{p}}_{n}}{2c}\big)-\sinh\big(\frac{{\bf{p}}_{n}}{2c}\big){\tilde{f}}^{-}({\bf{q}}_{n})\Big)=}\\ \\ \displaystyle{=4\sinh^{2}\big(\frac{{\bf{p}}_{n}}{2c}\big)\frac{\theta_{1}({\bf{q}}_{n}-{\bf{q}}_{n-1})\theta_{1}({\bf{q}}_{n}+{\bf{q}}_{n-1})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{n})}\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.}\end{array}$ (73)
Plugging (71), (72) and (73) into (69) we come to the following expression for $H$:
$\begin{array}[]{c}\displaystyle{H=(-1)^{n-1}\frac{2^{n}}{\theta_{1}^{\prime}(0)^{2n-2}}\,\frac{1}{\theta_{1}(2{\bf{q}}_{1})\theta_{1}(2{\bf{q}}_{n})}\,\left(\frac{\displaystyle{\prod\limits_{a=2}^{n}\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}}{\displaystyle{\prod\limits_{a=2}^{n-1}\theta_{1}(2{\bf{q}}_{a})}}\right)^{2}\,\prod\limits_{a=1}^{n}\sinh^{2}\big(\frac{{\bf{p}}_{a}}{2c}\big)\times}\\ \\ \displaystyle{\times\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.}\end{array}$ (74)
Next, let us once more use the identity (12):
$\begin{array}[]{c}\displaystyle{\wp({\bf{q}}_{a}-{\bf{q}}_{a-1})-\wp({\bf{q}}_{a}+{\bf{q}}_{a-1})=-\phi({\bf{q}}_{a}-{\bf{q}}_{a-1},{\bf{q}}_{a}+{\bf{q}}_{a-1})\phi({\bf{q}}_{a-1}-{\bf{q}}_{a},{\bf{q}}_{a}+{\bf{q}}_{a-1})=}\\ \\ \displaystyle{=-\frac{\theta_{1}^{\prime}(0)^{2}\theta_{1}(2{\bf{q}}_{a})\theta_{1}(2{\bf{q}}_{a-1})}{\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})^{2}\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})^{2}}\,.}\end{array}$ (75)
Then
$\begin{array}[]{c}\displaystyle{\prod\limits_{a=2}^{n}\frac{1}{\wp({\bf{q}}_{a}-{\bf{q}}_{a-1})-\wp({\bf{q}}_{a}+{\bf{q}}_{a-1})}=}\\ \\ \displaystyle{=\frac{(-1)^{n-1}}{\theta_{1}^{\prime}(0)^{2n-2}}\,\frac{1}{\theta_{1}(2{\bf{q}}_{1})\theta_{1}(2{\bf{q}}_{n})}\,\left(\frac{\displaystyle{\prod\limits_{a=2}^{n}\theta_{1}({\bf{q}}_{a}-{\bf{q}}_{a-1})\theta_{1}({\bf{q}}_{a}+{\bf{q}}_{a-1})}}{\displaystyle{\prod\limits_{a=2}^{n-1}\theta_{1}(2{\bf{q}}_{a})}}\right)^{2}\,.}\end{array}$ (76)
By comparing (74) and (76) one gets
$\begin{array}[]{c}\displaystyle{H=2^{n}\,\prod\limits_{a=2}^{n}\frac{1}{\wp({\bf{q}}_{a}-{\bf{q}}_{a-1})-\wp({\bf{q}}_{a}+{\bf{q}}_{a-1})}\,\prod\limits_{a=1}^{n}\sinh^{2}\big(\frac{{\bf{p}}_{a}}{2c}\big)\times}\\ \\ \displaystyle{\times\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.}\end{array}$ (77)
Finally, we come to the following Hamiltonian for the open elliptic Toda chain with boundaries:
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{openToda}}}=-\log(H)+\log 2^{n}=}\\ \\ \displaystyle{=\sum\limits_{a=1}^{n}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)-}\\ \\ \displaystyle{-\log\Big(\nu_{0}^{+}\coth\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)-\log\Big(\nu_{0}^{-}\coth\big(\frac{{\bf{p}}_{n}}{2c}\big)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)\,.\par}\end{array}$ (78)
To summarize, we proved the following statement.
Theorem The Hamiltonian for the open elliptic Toda chain with boundaries evaluated from the classical monodromy matrix ${\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (47) has the form (78), where the functions ${\tilde{f}}^{+}({\bf{q}}_{1})$, ${\tilde{f}}^{-}({\bf{q}}_{n})$ are defined in (64), (66) or (67). By construction, this model is gauge equivalent to the open XYZ chain with boundaries given by $K$-matrices (49), and with the Casimir functions fixed as in (34) at each site.
Compared to the original Hamiltonian (1) for the closed elliptic Toda chain, the expression (78) does not contain interaction of $n$-th particle with the first one. But it contains boundary terms for these two particles on boundaries.
@@
Example 1: pure open chain. Let $\nu^{\pm}_{k}=0$, $k=1,2,3$ and $\nu^{\pm}_{0}=1$. From viewpoint of the open XYZ chain this case corresponds to trivial $K$-matrices $K^{\pm}(z)=\sigma_{0}$ (49), and the gauge transformed $K$-matrices (52) are trivial as well: ${\tilde{K}}^{\pm}(z)=\sigma_{0}$. Then
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{pure\ open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)\cosh({\bf{p}}_{1}/2c)}+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)\cosh({\bf{p}}_{n}/2c)}+}\\ \\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (79)
This Hamiltonian differs from (1) by the absence of interaction between the first and the last particles, and by type of dependence on momenta for these two particles on the boundaries.
Example 2: only boundary terms. Now let it be vice versa: $\nu^{\pm}_{0}=0$. Using (67) we obtain
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{boundary}}}=\sum\limits_{a=1}^{n}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)-}\end{array}$ (80)
${-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\Big)-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp({\bf{q}}_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\Big)\,.}$
In this case the kinetic part (dependence on momenta) is the same as in (1) but there are additional external fields for the first and the last particles on the boundaries.
In conclusion, we notice again that the Ruijsenaars-Toda chain [2] generalizes the elliptic Toda chain (1), and it is gauge equivalent to a generic XYZ chain (without restrictions (34)). Similarly to the results of the present paper it is then possible to derive the open Ruijsenaars-Toda chain with boundaries. The boundary $K$-matrices can be also chosen to be dynamical, see the Zhukovsky-Volterra gyrostat [10] and the 4-constant ${\rm BC}_{1}$ Ruijsenaars-van Diejen model [11]. These generalizations will be studied elsewhere.
@@ Acknowledgments.
We are grateful to D. Murinov and A. Zabrodin for discussions.
This work was supported by the Russian Science Foundation under grant no. 25-11-00081,
https://rscf.ru/project/25-11-00081/ .